OpenStudy (anonymous):

Find the position versus time of a particle moving with the given acceleration data. a(t)=3cos(t)-2sin(t) s(0)=0, v(0)=4

OpenStudy (freckles):

seems to require integration

OpenStudy (freckles):

v=antiderivative of a use v(0)=4 to find integration constant s=antiderivative of v use s(0)=0 to find integration constant

OpenStudy (freckles):

do you know how to find the antiderivative of 3cos(t)-2sin(t) ?

OpenStudy (anonymous):

No not exactly

OpenStudy (freckles):

do you know derivative of sin(t) is cos(t) so the antiderivative of cos(t) is sin(t) do you know the derivative cos(t) is -sin(t) so the derivative of -cos(t) is sin(t) so the antiderivative of sin(t) is -cos(t)

OpenStudy (anonymous):

Yes I am just confused with the number part of it

OpenStudy (freckles):

antiderivative of 3*cos(t) is 3*sin(t) antiderivative of-2*sin(t) is -2*(-cos(t))

OpenStudy (freckles):

just bring the constant down

OpenStudy (freckles):

then take the antiderivative of the non-constant part

OpenStudy (anonymous):

Okay that makes more sense now. So when I am done with that what is the next step?

OpenStudy (freckles):

oops type-os

OpenStudy (anonymous):

And then after that I solve for D right?

OpenStudy (freckles):

\[a(t)=3\cos(t)-2\sin(t) \\ v(t)=\text{ antiderivative of } a(t) \\ v(t)=3\sin(t)-2(-\cos(t))+C \\ v(t)=3 \sin(t)+2 \cos(t)+C \\ \text{ we are given } v(0)=4 \\ v(0)=3\sin(0)+2\cos(0)+C \\ 4=3 \sin(0)+2\cos(0)+C \\ \text{ solve for C } \]

OpenStudy (freckles):

yes once we find this v(t) we need to do the antiderivative of v(t) to find s(t)

OpenStudy (freckles):

then use s(0)=0 to find your D

OpenStudy (anonymous):

Okay this makes so much more sense now thank you

OpenStudy (freckles):

np